OpenStudy (anonymous):

A car is driven along a level road . Total energy input from the petrol is 60kj & the car wastes 45kj of energy . What is the eficiency of the car?

OpenStudy (abhisar):

What do u think ?

OpenStudy (anonymous):

75%

OpenStudy (anonymous):

(45/60)*100

OpenStudy (abhisar):

Efficiency = \(\huge\frac{Output}{input}\)

OpenStudy (abhisar):

Here, car wastes 45KJ thus it gives an output of only 60-45=15KJ

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

yes

OpenStudy (abhisar):

*wasted

OpenStudy (abhisar):

so efficiency will be (15/60)*100=25%

OpenStudy (anonymous):

thanks dear friend

OpenStudy (abhisar):

\(\huge\color{red}\bigstar\)\(\huge\text{Rectification}\) \(\huge\color{red}\bigstar\) \(\color{blue}{\text{Originally Posted by}}\) @Abhisar Efficiency = \(\huge\frac{Output}{input}\) *100 \(\color{blue}{\text{End of Quote}}\)

OpenStudy (abhisar):

\(\color{red}{\huge\bigstar}\huge\text{You're Most Welcome! }\color{red}\bigstar\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)